Question 1. National Income Accounting (45 marks) Consider an economy where only three types of vehicles are produced and sold, namely: Tesla, GM and BMW. The following table shows the quantity sold (in millions) and the price per unit of each vehicle Tesla GM Quantity produced and sold 200 00 BMW 2011 40 $3 2012 40 S4. 2013 50 S4.50 210 S4. 014 2015 65 $5 300 Quantity produced and sold Price per unit 70 $5 375 $2 $3.75 12 S6 $2.50 $2.50 Price per unit Quantity produced and sold12 Price per unit 17 $7 20 $7 $4. $7 Copy and paste the above table in excel to conduct the computations for each question. Present your results clearly in a table as they apply to each question. a) Compute the yearly total value for each vehicle and the yearly nominal GDP for the economy. What is the growth rate in nominal GDP for each year? b) Assume 2012 is the base year. Compute the yearly real GDP for the economy. What is the growth rate in real GDP for each year? Compare this growth rate to the nominal GDP growth rate terms) in the overall price level relative to the base year? increase in the physical volume of output or the increase in prices? c) Compute the GDP deflator for each year. What is the yearly change (in percentage d) Would you say that the growth rate in nominal GDP for each year is due more to the 1 From 2011 to 2012, the price per unit of the three vehicles increased but the quantities did not. However, from 2014 to 2015, the quantities changed but the price per unit of each good did not change. In which case would you say that an average resident of this economy was better off? Explain briefly e)

Need help for (c), (d) and (e)
Answer #1

To find the solution first of all have to find out the nominal GDP and real GDP. Nominal GDP is the value of goods and services estimated at the current year prices. In other words Nominal GDP is known as the GDP at current prices. It is calculated by multiplying the current year quantity with the current year prices.   But the value may be increase due to the inflation. Such value addition is not excluded in Nominal GDP. The formula for calculating Nominal GDP is Nominal GDP =Q × Prices (current year) the current year means the price prevailing in the year of manufacturing.

Nominal GDP

2011

2012

2013

2014

2015

Product

Q

P

GDP

Q

P

GDP

Q

P

GDP

Q

P

GDP

Q

P

GDP

Tesla

40

3

120

40

4

160

50

4.50

225

65

5

325

70

5

350

GM

200

2

400

200

3.8

750

210

4

840

300

2.50

750

375

2.5

938

BMW

12

4

48

12

6

72

13

7

91

17

7

119

20

7

140

Total

568

982

1156

1194

1428

Real GDP is the real index of economic growth. It is the money value of all goods and services produced within the domestic territory of a country in one year after adjusting for inflation. For adjusting inflation we choose one year as base year and GDP is calculated in terms of the prices in that year. In following such method we can find out the GDP growth in actual and exact value since the value addition from inflation is excluded. Real GDP is otherwise known as GDP at constant price. Real GDP = Q × Base year price. Base year price means the year of price chosen for comparison.

Real GDP (2011 is taken as the base year, multiply current year quantity with base year price)

2011

2012

2013

2014

2015

Product

Q

P

GDP

Q

P

GDP

Q

P

GDP

Q

P

GDP

Q

P

GDP

Tesla

40

3

120

40

3

120

50

3

150

65

3

195

70

3

210

GM

200

2

400

200

2

400

210

2

420

300

2

600

375

2

750

BMW

12

4

48

12

4

48

13

4

52

17

4

68

20

4

80

Total

568

568

622

863

1040

c. GDP deflator

GDP deflator measures the quantum of inflation on the GDP during a year. The GDP deflator compares the Nominal GDP and Real GDP and find out the impact of inflation on the GDP. In the absence of inflation there is no difference between the nominal GDP and Real GDP. In such situation the Nominal GDP itself is the Real GDP.

The formula for finding GDP deflator is Nominal GDP/Real GDP× 100

GDP deflator for the year 2012 is 982/568×100=172.9

GDP deflator for the year 2013 is 1156/622×100=185.85

GDP deflator for the year 2014 is 1194/863×100= 138.36

GDP deflator for the year 2015 is 1428/1040×100=137.30

The inflation or the increase in price level can be found out by using the formula GDP deflator of the current year - the GDP deflator of the previous year/GDP deflator of the previous year×100.

i.e    GDP2-GDP1

           _______                   × 100.

             GDP1

The inflation rate for the year 2012 is 172.9- 100/100×100= 72

The inflation rate for the year 2013 is 185.85-172.9/172.9×100= 7.48

The inflation rate for the year 2014 is 138.36-185.85/185.85×100=-25

The GDP deflator for the year 2015 is 137.30-138.36/138.36×100=-0.76

d. The nominal GDP will be exaggerated due to inflation. The Nominal GDP shows an increase in the value of GDP due to inflation without an increase in the physical volume of output. But if the price level is constant, the Nominal GDP increases with the increase in the physical volume of output. In the years 2012 and 2013 the Nominal GDP increases with inflation. But in the years 2014 and 2015 the Nominal GDP increased due to the increase in physical volume of output. During these years the inflation rate was minus.

e. The citizens of a country is better off if there is an increase in Real GDP. The increase in Real GDP occurs when there is an increase in the quantity of output without inflation. If the value of GDP increase with price increase, the citizen will not be better off. In the years 2011 and 2012 the GDP shows an increase simply due to an increase in price level. Then the citizens are not better off. But in the period from 2014 to 2015 GDP increase with the increase in the volume of output at constant price. Hence in these periods the citizens are better off than before.